# TBAC-UniImage: Unified Understanding and Generation by Ladder-Side Diffusion Tuning

### v1

@@ 4 Conclusion
This paper introduces TBAC-UniImage, a novel architecture designed to unify multimodal understanding and generation. Our approach proposes a direct link between a Multimodal Large Language Model (MLLM), which excels at understanding, and a Diffusion Transformer (DiT), a powerful generative model. The core idea of integration is our proposed Ladder-Side Diffusion Tuning mechanism. This technique treats the DiT as a generative ladder for the MLLM, channeling hidden states from various MLLM layers to the DiT. These states serve as rich, multi-level conditional inputs, guiding the image generation process with nuanced understanding. Extensive evaluations on three widely-used benchmarks demonstrate the efficacy of our method. Nevertheless, we acknowledge several limitations that present avenues for future work: (1) enhancing the model's fine-grained comprehension of complex, dense prompts; (2) improving generation consistency for image-text-to-image tasks; and (3) refining the in-image text rendering capability. In conclusion, we posit that TBAC-UniImage and our Ladder-Side Diffusion Tuning paradigm offer a valuable new perspective for the research and development of next-generation unified models.

### v2

@@ 4 Conclusion
This paper introduces TBAC-UniImage, a novel architecture designed to unify multimodal understanding and generation. Our approach proposes a direct link between a Multimodal Large Language Model (MLLM), which excels at understanding, and a Diffusion Transformer (DiT), a powerful generative model. The core idea of integration is our proposed Ladder-Side Diffusion Tuning mechanism. This technique treats the DiT as a generative ladder for the MLLM, channeling hidden states from various MLLM layers to the DiT. These states serve as rich, multi-level conditional inputs, guiding the image generation process with nuanced understanding. Extensive evaluations on four widely-used benchmarks demonstrate the efficacy of our method. Nevertheless, we acknowledge several limitations that present avenues for future work: (1) enhancing the model's fine-grained comprehension of complex, dense prompts; (2) improving generation consistency for image editing tasks; and (3) refining the in-image text rendering capability. In conclusion, we posit that TBAC-UniImage and our Ladder-Side Diffusion Tuning paradigm offer a valuable new perspective for the research and development of next-generation unified models.